# Decentralization of Ethereum’s Builder Market

### v3

@@ 4.1 Competitiveness of MEV-Boost Auctions
Ideally, revenue from MEV-Boost auctions should be near optimal due to competition. However, whether real-world auctions induce such near-perfect competition is complicated by several confounding factors: builders may not be able to respond to bids due to network latency, or they may collude implicitly or explicitly. In this section, we quantify the competitiveness of the MEV-Boost auctions using what we call competitive index (CI).
We also look at the allocation efficiency as it evaluates whether the current mechanism induces the desired outcomes.
@@
To compute $TV(B)$ of a given block $B$, we follow the standard approach in open source MEV analysis tools [20, 31] and previous works [41, 80], though we built our own analysis toolchain and optimized its performance to handle the large volume of auction data. We compute $EV(B)$ using a modified Ethereum execution client (Reth [68]) to execute transactions in $B$ and compute the net ETH balance changes of the bidder’s addresses. Most builders’ addresses are well-known partly because they need to build up a reputation. We cross-validated it with other sources to ensure that our dataset was the most complete one to date (see Sec. 3.1 for details.) Then, with bid value $BV(B)$ from the auction dataset, the true value $TV(B)$ is the sum of $EV(B)$ and $BV(B)$.
In every slot $s$, bidders can submit multiple bids throughout the duration of the slot. We use $BV(p,s)$ to denote the highest bid value from bidder $p$ in slot $s$, and $TV(p,s)$ the corresponding true value.
@@
Namely, $CI(s)$ measures the relative difference between the winning bid and the second-highest true value. $CI(s)\geq 0$ indicates that the winning bid value is not less than the second-highest true value, satisfying the definition of a competitive auction, while $CI(s)<0$ means insufficient competition.
@@
Analysis of uncompetitiveness and inefficiency.   Fig. 4 shows the calculated CI and EI distribution across all MEV-Boost auctions in the ultra sound relay dataset. Notably, 91.8% (sum of the areas with $CI\geq 0$) of the auctions are competitive, yet only 75.1% (sum of the areas with $EI>0$) are efficient.
@@
In the interest of space, we leave the analysis of efficiency and the implication of MEV on efficiency to Appx. D.1.
MEV’s implication on auction outcomes.   To further understand how the magnitude of MEV in a slot affects the competitiveness and efficiency of the auction, we categorize every slot into one of three tiers according to the winning bid (as a proxy to the significance of the MEV opportunity in that slot): low MEV (below 0.0215 ETH, the 10th percentile of the winning bid values), high MEV (above 0.3933 ETH, 95th percentile) and medium MEV in between. The ratio of competitive auctions in each tier is shown in Fig. 5. We note that competitiveness decreases as the MEV of a slot increases.
@@
Loss from uncompetiveness.   For an uncompetitive auction, the loss from uncompetitiveness is the difference between the second-highest true value and the actual winning bid value, representing the loss a validator incurs compared to what it would have received in a competitive auction. We computed the loss from uncompetitiveness using our dataset, and the results are shown in Tab. 2. The loss from uncompetitiveness ranges from 0.5% to 1.7% during the periods covered.
Table 2: Losses from uncompetitiveness and inequality.
@@ 4.2 Inequality in Block-building Capacities
QCD is a real value in $[0,1]$. A larger QCD indicates greater dispersion (the intuition is that in a dataset with big dispersion, $Q_{3}$ is much larger than $Q_{1}$, thus $QCD$ approaches $1$). Since it uses quartiles, QCD is robust to outliers [88].
@@
In Fig. 6, we plot the distribution of the QCD for MEV-Boost auctions from April 2023 to December 2023 for different builder groups and MEV tiers. (We sample the first week of each month. See Sec. 3.1 for details.) It reveals two trends. First, within each MEV tier, there is a clear increase in inequality (a high QCD means high inequality) from the top to the middle and then to the tail. This suggests that while top builders have comparable block-building capacities, middle and tail builders have significantly different capacitiesx. This might be good news: top builders having similar abilities means they can meaningfully compete, which is positive for the decentralization of the builder market. Second, the inequality worsens as the MEV tier increases. One possible reason is that providers may give significant MEV opportunities to select builders, as we will discuss in Sec. 5.3.
@@
As shown in Tab. 2, the loss from inequality ranges from 5.6% to 11.5% during the periods covered by our dataset. Comparing the two types of loss, we find that inequality is the proposers’ primary source of loss. Note that the time range of the full bids is from April to December 2023, and the inequality in the current builder market may worsen (as discussed in Sec. 5.3).
5 What Caused Inequality in Block-building Capacities
Given the significant losses proposers incur due to the inequality in builders’ block-building capacity, we further investigate the causes of this inequality in this section.
Although the builder market is technically open to all, the barrier to accessing profitable private order flows (Sec. 2.2) forms an implicit access barrier.
To illustrate the importance of private order flow to the success of builders, Fig. 7 plots the fraction of builder income from private order flows. Currently, approximately 80% of the block value comes from private order flows.
Figure 7: Fraction of builder income from private order flows.
Finding III: The private order flow is crucial for builders to win the MEV-Boost auction because it contributes to more than 80% of the block value in over 50% of blocks.
Where can new builders get private order flows? The entities who provide (or sell) private order flows to builders are called private order flow providers. Different providers impose different access barriers to their private order flows. Some providers, like MEV Blocker or MEV-Share, are open to new builders who satisfy certain requirements, while some (e.g., most searchers) are almost opaque to the public.
To understand what causes the access barrier for new builders, we first identify important “pivotal” providers and then analyze the requirements to access private order flows from these representative pivotal providers.
@@ 5.1 Pivotal Private Order Flow Providers
Prominent private order flow providers are listed in Tab. 5. Based on their roles in the MEV supply chain (see Fig. 3), we categorize them into three types: searcher, Telegram trading bot, and channel.
We now present our method to quantify the importance of individual providers. An existing approach [53, 54] is to compute the contribution of each provider to the builders’ income (Sec. 2.3) during a specific period. However, this approach only considers the magnitude of impact but neglects the duration (in terms of time or the number of auctions). For example, a single private transaction, such as one3330xf0464b01d962f714eee9d4392b2494524d0e10ce3eb3723873afd1346b8b06e4 transferring 457.55 ETH to the builder as a tip, could contribute significantly to the builder’s income, but this provider may only influence the MEV-Boost auction in a single slot. Thus, we assess the significance of a provider as the number of auctions for which they played an important role; in particular, we focus on the providers that can affect half of the MEV-Boost auctions in over two weeks, thereby capturing their sustained impact.
@@
Figure 8: A provider $P$ is pivotal if the winner of an auction could have lost without order flows from $P$. The pivotal level of top-5 pivotal providers over time. A line may not start from the beginning because the provider did not exist at that point. Maestro became an independent provider around Dec 2023.

### v4

@@ 4.1 Competitiveness of MEV-Boost Auctions
Ideally, revenue from MEV-Boost auctions should be near optimal due to competition. However, whether real-world auctions induce such near-perfect competition is complicated by several confounding factors: builders may not be able to respond to bids due to network latency, or they may collude implicitly or explicitly. In this section, we quantify the competitiveness of the MEV-Boost auctions using what we call competitive index (CI). We also look at the allocation efficiency as it evaluates whether the current mechanism induces the desired outcomes.
@@
To compute $TV(B)$ of a given block $B$, we follow the standard approach in MEV analysis tools [41, 42] and previous works [43, 44], though we built our own analysis toolchain and optimized its performance to handle the large volume of historical auction data (326.6 million full bids). We compute $EV(B)$ using a modified Ethereum execution client (Reth [47]) to execute transactions in $B$ and compute the net ETH balance changes of the builder’s addresses. Most builders’ addresses are well-known partly because they need to build up a reputation. We cross-validated it with other sources to ensure that our dataset was the most complete one to date (see Sec. 3.1 for details.) Then, with bid value $BV(B)$ from the auction dataset, the true value $TV(B)$ is the sum of $EV(B)$ and $BV(B)$.
In every slot $s$, builders can submit multiple bids throughout the duration of the slot. We use $BV(p,s)$ to denote the highest bid value from builder $p$ in slot $s$, and $TV(p,s)$ the corresponding true value.
@@
Namely, $CI(s)$ measures the relative difference between the winning bid and the second-highest true value. $CI(s)\geq 0$ indicates that the winning bid value is not less than the second-highest true value, thus satisfying the definition of a competitive auction, while $CI(s)<0$ means there is insufficient competition.
@@
Analysis of uncompetitiveness and inefficiency.   Fig. 4 shows the calculated CI and EI distribution across all MEV-Boost auctions in the our dataset. Notably, 91.8% (sum of the areas with $CI\geq 0$) of the auctions are competitive, yet only 75.1% (sum of the areas with $EI>0$) are efficient.
@@
MEV’s implication on auction outcomes.   To further understand how the magnitude of MEV in a slot affects the competitiveness and efficiency of the auction, we categorize every slot into one of three tiers based on the winning bid (as a proxy for MEV significance in that slot): low MEV (below 0.0215 ETH, the 10th percentile of the winning bid values), high MEV (above 0.3933 ETH, 95th percentile) and medium MEV in between. The ratio of competitive auctions in each tier is shown in Fig. 5. We note that competitiveness decreases as the MEV of a slot increases.
@@
In the interest of space, we leave the analysis of efficiency to Appx. D.1.
Loss from uncompetiveness.   For an uncompetitive auction, the loss from uncompetitiveness is the difference between the second-highest true value and the actual winning bid value, representing the loss a validator incurs compared to what it would have received in a competitive auction. We computed this loss using our dataset, and the results are shown in Tab. II. The loss from uncompetitiveness ranges from 0.5% to 1.7% during our study periods.
TABLE II: Losses from uncompetitiveness and inequality.
@@ 4.2 Inequality in Block-building Capacities
QCD is a real value in $[0,1]$. A larger QCD indicates greater dispersion (the intuition is that in a dataset with big dispersion, $Q_{3}$ is much larger than $Q_{1}$, thus $QCD$ approaches $1$). Since QCD uses quartiles, this metric is robust to outliers [16].
@@
In Fig. 6, we plot the distribution of the QCD for MEV-Boost auctions from April 2023 to December 2023 for different builder groups and MEV tiers. (We sample the first week of each month. See Sec. 3.1 for details.) It reveals two trends. First, within each MEV tier, there is a clear increase in inequality (a high QCD means high inequality) from the top to the middle and then to the tail. This suggests that while top builders have comparable block-building capacities, middle and tail builders have significantly different capacities. This might be good news: top builders having similar abilities means they can meaningfully compete, which is positive for the decentralization of the builder market. Second, the inequality worsens as the MEV tier increases. One possible reason is that providers may give significant MEV opportunities to select builders, as we will discuss in the next section.
@@
As shown in Tab. II, the loss from inequality ranges from 5.6% to 11.5% during the periods covered by our dataset. Comparing the two types of loss, we find that inequality is the primary source of proposers’ loss.
4.3 Proposer Losses Over Extended Periods
Note that the time range of the full bids is from April to December 2023, raising the natural question of whether proposers may incur more significant losses in 2024.
Recall that, in an MEV-Boost auction in the slot $s$, the proposer’s loss is calculated as $TV(p_{1},s)-BV_{w}(s)$. Unfortunately, without the full bid dataset, we cannot compute the highest true value $TV(p_{1},s)$ and, therefore, the proposer loss. However, since the winner’s true value cannot exceed the highest true value (otherwise, it would be the highest itself), we can compute a lower bound for the actual proposer losses using $TV_{w}(s)-BV_{w}(s)$, which we refer to as the estimated proposer loss. A significant estimated proposer loss suggests that proposers incur substantial losses.
We calculate the daily estimated proposer loss by aggregating the estimated losses from all MEV-Boost auctions each day and comparing this total to the overall block value for that day. As shown in Fig. 7, there is a clear upward trend in the daily estimated proposer loss in 2024—the median was 4.1% before, compared to 21.2% afterward! The daily estimated proposer loss is significant on August 6, 2024, primarily because, in an auction on that date (slot 9676257), the winner’s bid is 11 ETH while the true value is 803 ETH—resulting in a significant loss for the proposer.
Figure 7: Daily estimated proposer loss over two years. Proposers suffered significant losses in 2024.
Why are proposer losses more significant in 2024? One possible explanation is that worsening inequality in 2024 amplifies proposer losses. To explore this, we need to deepen our understanding of inequality in the builder market.
5 Private Order Flows
In the previous section, our study on historical auctions reveals that inequality is a primary source of proposers’ losses. Recall that the inequality loss comes from the disparity in builders’ true values—a bid’s true value is computed by the total value extracted from the transactions. As introduced in Sec. 2.2, transactions in Ethereum have two ways to reach a builder: through the public peer-to-peer network or via a direct channel. The former is publicly available to all, which is less concerning, while the latter is restricted, providing access only to select builders.
In this section, we further study the critical factor contributing to inequality—the difference in private order flows that builders receive.
Figure 8: Fraction of block value from private order flows.
Criticalness of private order flows.   To illustrate the importance of private order flow to builders, we compute the ratio of block values derived from private transactions. We first identify all private order flows using data described in Sec. 3.1. Then, for each block, we compute the ratio between values extracted from private transactions (i.e., the total priority fees [45] and direct transfers [46] to the block builder) to the block value. Fig. 8 plots the fraction of block value from private order flows over two years. As of October 2024, approximately 80% of the block value comes from private order flows for the majority of the blocks, suggesting the criticalness of private order flows!
Finding III: Private order flows are critical to builders’ block-building capacities since they contribute to more than 80% of the block value in over 50% of blocks.
@@ 5.1 Pivotal Private Order Flow Providers
Private order flows are critical to builders, and they can drive significant differences in builders’ block-building capacities, contributing to inequality. Therefore, analyzing them is essential for understanding such inequality, despite the challenges posed by the vast data volume and extended time span (112 million private transactions over two years).
Our first step is to classify the large volume of private order flows based on their sources. The entities who provide (or sell) private order flows to builders are called private order flow providers, or providers for short. Prominent private order flow providers are listed in Tab. V. Based on their roles in the MEV supply chain (see Fig. 3), we categorize them into three types: searcher, Telegram trading bot, and channel. The process of identifying the order flows from these providers, as well as the details of the finalized dataset, are described in Sec. 3.1.
Not every provider is equally important. Some providers generate more value than others and their order flows can significantly impact builders. Therefore, our next step is to identify these important providers.
An existing approach [18, 19] is to compute the contribution of each provider to the builders’ income (Sec. 2.3) during a specific period. However, this approach can have false positives; for example, it could identify providers that were only influential temporarily. For example, a single private transaction3330xf0464b01d962f714eee9d4392b2494524d0e10ce3eb3723873afd1346b8b06e4 transferring 457.55 ETH to the builder as a tip could contribute significantly to the builder’s income, but this provider may only influence the MEV-Boost auction in a single slot. Thus, we assess the significance of a provider as the number of auctions for which they played an important role; in particular, we focus on the providers that have sustained impacts (e.g., they can affect half of the MEV-Boost auctions in over two weeks).